# Hungarian Research Network (HUN-REN) Laboratory on Engineering and Management Intelligence Research Group of Operations Research and Decision Systems

### v12

@@ Abstract
Several sports tournaments contain a round-robin group stage where the teams are assigned to groups subject to some constraints. Hence, the organisers usually use a computer-assisted random draw to avoid any dead end, a situation when the teams still to be drawn cannot be assigned to the remaining empty slots. This procedure is known to be unfair: the feasible allocations are not equally likely. We quantify the unfairness of the 2018 FIFA World Cup draw and evaluate its effect on the probability of qualification for the knockout stage for each national team. The official draw order of Pot 1, Pot 2, Pot 3, Pot 4 turns out to be a significantly better option than the 23 other draw orders with respect to the unwanted distortions. Nonetheless, the non-uniform draw distorts the probability of qualification by more than one percentage point for two countries. Our results call attention to the non-negligible role of draw order and make it possible for policymakers to decide whether the use of fairer draw mechanisms is justified.
@@ 1 Introduction
However, this mechanism, based on a computer-assisted random draw to avoid any dead end (a situation when the teams still to be drawn cannot be assigned to the remaining empty slots), is not uniformly distributed: the valid assignments are not equally likely to occur (Roberts and Rosenthal,, 2023). Hence, Roberts and Rosenthal, (2023) propose two unbiased mechanisms that use balls and bowls, making them suitable for a televised draw. On the other hand, they require random simulations, which might threaten transparency.
@@
The current work aims to fill this research gap. In particular, the 2018 FIFA World Cup has been chosen as a case study since the 2022 FIFA World Cup draw has contained an inherent bias due to the uncertainty in the set of teams qualified (Csató,, 2023), and the teams qualifying for the 2026 FIFA World Cup are naturally unknown at the time of writing.
We also call attention to the role of draw order: the order of the pots in the group draw turns out to have a non-negligible sporting effect. This is a somewhat surprising finding because the fairness of the UEFA Champions League Round of 16 draw essentially does not depend on whether the group winners or the runners-up are drawn first (Boczoń and Wilson,, 2023; Klößner and Becker,, 2013).
Our topic is hugely relevant to governing bodies in sports. In an ideal environment, the rules governing a sport prevent any player or team from gaining an unfair advantage (Csató,, 2021; Devriesere et al.,, 2024). But the sequential draw method of FIFA may threaten fairness, while the alternative solutions remain more complex and less transparent (Roberts and Rosenthal,, 2023). Therefore, policymakers could maintain the balance between fairness and integrity only by investigating different fairness indicators (Scelles et al.,, 2024).
@@ 2 Related literature
Laliena and López, (2019) develop two uniformly distributed designs for group draw with geographical restrictions that produce groups having similar or equal competitive levels. Cea et al., (2020) analyse the deficiencies of the 2014 FIFA World Cup draw and provide a mixed integer linear programming model to create the groups. The suggested method takes into account draw restrictions and aims to balance “quality” across the groups.
@@ 3 Theoretical background
A permutation of an ordered set is a rearrangement of its elements. In the FIFA World Cup draw, the initial order of the teams is provided by a random draw. In an unrestricted group draw, the teams can be assigned to the groups in this order. However, in the presence of draw conditions, it is not obvious to find the permutation of the teams that corresponds to the feasible allocation implied by the draw procedure of FIFA.
@@
Assume that there are $k=4$ groups A–D and $n=4$ teams $T1$$T4$ drawn sequentially from a pot. The order of permutations to be checked according to the draw mechanism is shown in Figure 1. Team $T1$ is assigned to group A in the first six permutations because it can be placed in another group only if either group A is unavailable for team $T1$ or teams $T2$$T4$ cannot be assigned to groups B–D.
@@
Generating all permutations of a sequence of values in a specific order is a famous problem in computer science (Sedgewick,, 1977). The classic algorithm of lexicographic ordering goes back to Narayaṇa Paṇḍita, an Indian mathematician from the 14th century (Knuth,, 2005). The ordering corresponding to the FIFA World Cup draw procedure is called representation via swaps (Arndt,, 2010) and has been presented first in Myrvold and Ruskey, (2001) according to our knowledge. Arndt, (2010, Figure 10.1-E) contains the same order of permutations as Figure 1.
@@
The description of the 2018 FIFA World Cup draw (FIFA,, 2017) does not give an exact algorithm to obtain the implied feasible allocation of the teams into groups for a given random order. The scheme of an appropriate computer program is presented in Figure 2. Pot $j$, from which team $i$ is drawn, is not a variable, however, the number of available groups can be different for each pot. The algorithm is based on backtracking: if the remaining teams cannot be assigned to the empty group slots in any order such that all restrictions are satisfied, then the last team is placed in the next available group in alphabetical order. This process is repeated until the associated feasible allocation is obtained or the non-existence of a valid assignment is verified.
Backtracking can be familiar from the problem of scheduling round-robin tournaments, where an unlucky assignment of games to slots can result in a schedule that cannot be completed (Rosa and Wallis,, 1982; Schaerf,, 1999). Backtracking is also widely used to solve puzzles such as the eight queens puzzle, crosswords, or Sudoku.
@@ 24 Iran AFC 1790 32 Saudi Arabia AFC 1586
4.1 The effect of draw order
According to Roberts and Rosenthal, (2023, Section 2), the FIFA World Cup draw procedure is unfair since some feasible allocations might occur with a higher probability. In addition, the pre-assignment of Russia to Group A introduces a powerful bias because the draw mechanism is not independent of group labels. Russia has a 12.5% probability of playing against an arbitrarily chosen country from Pot 2 because no draw constraints can apply. However, since there are one CONCACAF, three CONMEBOL, and four UEFA members in Pot 2, the two CONMEBOL teams from Pot 1 (Brazil and Argentina) play against a given European team from Pot 2 with a probability of 0.2 since they have five possible opponents from Pot 2. The remaining five UEFA teams in Pot 1 are identical concerning the draw constraints, thus, they have a chance of $(100-12.5)/5=17.5$% to be assigned to the same group as a given South American team from Pot 2. Analogously, the probability that two given European teams from Pots 1 and 2 play against each other equals $(100-2\times 20-12.5)/5=9.5$%.
However, an appropriate relabelling of the pots may bring the FIFA World Cup draw closer to the uniform distribution. Therefore, all possible orders of the pots are examined such that the pre-assignment of Russia to Group A is retained. Since the teams can be drawn in $7!\times\left(8!\right)^{3}\approx 3.3\times 10^{17}$ different orders, it is almost impossible to derive exact theoretical results. Consequently, the 25 draw mechanisms—the draw procedure of FIFA with the 24 possible orders of the pots and the uniformly distributed rejection mechanism (Roberts and Rosenthal,, 2023)—are analysed on the basis of 1 million randomly generated draw orders, analogous to Roberts and Rosenthal, (2023).
@@ 4.1 The effect of draw order
where $p_{ij}^{M}$ and $p_{ij}^{U}$ are the probabilities that teams $i$ and $j$ are assigned to the same group by a mechanism $M$ and the uniform draw, respectively. When the biases are aggregated for the teams, we will always sum their absolute value $\lvert\Delta_{ij}\rvert$.
@@
The distortions are also worth studying from the perspective of the countries by taking the average of absolute or squared biases for the pairs that involve the given national team. For the sake of simplicity, four draw orders will be analysed in detail: the official 1-2-3-4, one that improves both the minimal nonzero and maximal probabilities with a slight change (1-2-4-3), one that is less biased for Mexico (1-3-2-4), and a fundamentally different order (4-2-3-1).
@@
Figure 3 compares the means of the biases for the 32 countries. The official draw procedure is far from the fair rejection mechanism in the case of Russia (1), Brazil (3), Argentina (5), Mexico (14), and Serbia (25). Note that Brazil and Argentina are interchangeable in the draw, thus, the deviations for these two countries are different only due to the (seemingly low) sampling error. Reversing the order of Pots 3 and 4 reduces unfairness for Russia and Serbia but it does not treat the problem for Mexico. Contrarily, the deviation for Costa Rica (19) strongly increases with the draw order 1-2-4-3. Drawing Pot 3 immediately after the best teams in Pot 1 implies a relatively low distortion in the case of Russia, and, especially, Mexico but the situation of the African (21, 22, 23) and European (17, 18, 20) teams in Pot 3 becomes less fair compared to the draw order 1-2-4-3. Finally, the draw order 4-2-3-1 is unfair with respect to all countries in Pot 2 except for Mexico.

### v13

@@ Abstract
Several sports tournaments contain a round-robin group stage where the teams are assigned to groups subject to some constraints. Hence, the organisers usually use a computer-assisted random draw to avoid any dead end, a situation when the teams still to be drawn cannot be assigned to the remaining empty slots. This procedure is known to be unfair: the feasible allocations are not equally likely, that is, the draw does not have a uniform distribution. We quantify the implied unfairness of the 2018 FIFA World Cup draw and evaluate its effect on the probability of qualification for the knockout stage for each national team. The official draw order of Pot 1, Pot 2, Pot 3, Pot 4 turns out to be a significantly better option than the 23 other draw orders with respect to the unwanted distortions. Nonetheless, the non-uniform draw distorts the probability of qualification by more than one percentage point for two countries. Our results call attention to the non-negligible role of draw order and make it possible for policymakers to decide whether using fairer draw mechanisms is justified.
@@ 1 Introduction
However, this mechanism, based on a computer-assisted random draw to avoid any dead end (a situation when the teams still to be drawn cannot be assigned to the remaining empty slots), is not uniformly distributed: the valid assignments are not equally likely to occur (Roberts and Rosenthal,, 2024). Hence, Roberts and Rosenthal, (2024) propose two unbiased mechanisms that use balls and bowls, making them suitable for a televised draw. On the other hand, they require random simulations, which might threaten transparency.
@@
The current work aims to fill this research gap by analysing the unique mechanism used in practice by sports federations to draw groups (containing at least three teams) with constraints through the example of the 2018 FIFA World Cup. This case study has been chosen since the 2022 FIFA World Cup draw has contained an inherent bias due to the uncertainty in the set of teams qualified (Csató,, 2023), and the teams qualifying for the 2026 FIFA World Cup are naturally unknown at the time of writing.
We also call attention to the role of draw order: the order of the pots in the group draw turns out to have a non-negligible sporting effect. This is a somewhat surprising finding because the fairness of the UEFA Champions League Round of 16 draw essentially does not depend on whether the group winners or the runners-up are drawn first (Boczoń and Wilson,, 2023; Klößner and Becker,, 2013). On the other hand, no clear recommendation is given for the optimal draw order in general due to the huge complexity of the calculations; according to Csató, (2024, Section 5.1), computing the probabilities for six teams drawn from three pots is barely possible—and the 2018 FIFA World Cup contains 32 teams and eight pots. Therefore, several more case studies would be needed to find a general pattern if it exists.
Our topic is hugely relevant to governing bodies in sports. In an ideal environment, the rules governing a sport prevent any player or team from gaining an unfair advantage (Csató,, 2021; Devriesere et al.,, 2024). But the sequential draw method of FIFA may threaten fairness, while the alternative solutions remain more complex and less transparent (Roberts and Rosenthal,, 2024). Therefore, policymakers could maintain the balance between fairness and integrity only by investigating different fairness indicators (Scelles et al.,, 2024).
@@ 2 Related literature
Laliena and López, (2019) develop two uniformly distributed designs for group draw with geographical restrictions that produce groups having similar or equal competitive levels. Cea et al., (2020) analyse the deficiencies of the 2014 FIFA World Cup draw and provide a mixed integer linear programming model to create the groups. The suggested method takes draw restrictions into account and aims to balance “quality” across the groups.
@@ 3 Theoretical background
A permutation of a set is a rearrangement of its elements. In the FIFA World Cup draw, the initial permutation of the teams is provided by a random draw. In an unrestricted group draw, the teams can be assigned to the groups in this permutation. However, in the presence of draw conditions, it is not obvious to find the permutation of the teams that corresponds to the feasible allocation implied by the draw procedure of FIFA.
@@
Assume that there are $k=4$ groups A–D and $n=4$ teams $T1$$T4$ drawn sequentially from a pot. The sequence of permutations to be checked according to the draw mechanism is shown in Figure 1. Team $T1$ is assigned to group A in the first six permutations because it can be placed in another group only if either group A is unavailable for team $T1$ or teams $T2$$T4$ cannot be assigned to groups B–D.
@@
    The draw procedure of FIFA works as follows:
    • $\diamond$
      Team $T1$ cannot be assigned to the first empty slot in group A, hence, it is placed in the first available group, which is group B.
    • $\diamond$
      Team $T2$ is assigned to the first empty slot in group A.
    • $\diamond$
      Team $T3$ cannot be assigned to the first empty slot in group C, hence, it is placed in the first available group, which is group D.
    • $\diamond$
      Team $T4$ is assigned to the first (and last) empty slot in group C.
@@
    The draw procedure of FIFA works as follows:
    • $\diamond$
      No constraint prohibits directly the assignment of team $T1$ to group A. However, if team $T1$ is not assigned to group C, then the three remaining teams $T2$$T4$ cannot be assigned to groups B–D. A similar argument uncovers that team $T1$ should be assigned to group C.
    • $\diamond$
      Team $T2$ cannot be assigned to the first empty slot in group A, hence, it is placed in the first available group, which is group B.
    • $\diamond$
      Team $T3$ is assigned to the first empty slot in group A.
    • $\diamond$
      Team $T4$ is assigned to the first (and last) empty slot in group D.
The official video of the 2018 FIFA World Cup group draw is available at https://www.youtube.com/watch?v=jDkn83FwioA.
Generating all permutations of a sequence of values is a famous problem in computer science (Sedgewick,, 1977). The classic lexicographic algorithm goes back to Narayaṇa Paṇḍita, an Indian mathematician from the 14th century (Knuth,, 2005). The sequence corresponding to the FIFA World Cup draw procedure is called representation via swaps (Arndt,, 2010) and has been presented first in Myrvold and Ruskey, (2001) according to our knowledge. In particular, Figure 1 is analogous to Arndt, (2010, Figure 10.1-E).
@@
The description of the 2018 FIFA World Cup draw (FIFA,, 2017) does not give an exact algorithm to obtain the implied feasible allocation of the teams into groups for a given random permutation of the teams. The scheme of an appropriate computer program is presented in Figure 2. Pot $j$, from which team $i$ is drawn, is not a variable, however, the number of available groups can be different for each pot. The algorithm is based on backtracking: if the remaining teams cannot be assigned to the empty group slots in any permutation such that all restrictions are satisfied, then the last team is placed in the next available group in alphabetical order. This process is repeated until the associated feasible allocation is obtained or the non-existence of a valid assignment is verified.
Backtracking can be familiar from the problem of scheduling round-robin tournaments, where an unlucky assignment of games to slots can result in a schedule that cannot be completed (Rosa and Wallis,, 1982; Schaerf,, 1999). Backtracking is also widely used to solve puzzles such as the eight queens puzzle, crosswords, or Sudoku. For draw procedures under draw constraints, backtracking has been suggested first in Guyon, (2014).
@@ 24 Iran AFC 1790 32 Saudi Arabia AFC 1586
4.1 The effect of the draw order
According to Roberts and Rosenthal, (2024, Section 2), the FIFA World Cup draw procedure is unfair since some feasible allocations might occur with a higher probability. In addition, the pre-assignment of Russia to Group A introduces a powerful bias because the draw mechanism is not independent of group labels. Russia has a 12.5% probability of playing against an arbitrarily chosen country from Pot 2 because no draw constraints can apply. However, since there are one CONCACAF, three CONMEBOL, and four UEFA members in Pot 2, the two CONMEBOL teams from Pot 1 (Brazil and Argentina) play against a given European team from Pot 2 with a probability of 0.2 since they have five possible opponents from Pot 2. The remaining five UEFA teams in Pot 1 are identical concerning the draw constraints, thus, they have a chance of $(100-12.5)/5=17.5$% to be assigned to the same group as a given South American team from Pot 2. Analogously, the probability that two given European teams from Pots 1 and 2 play against each other equals $(100-2\times 20-12.5)/5=9.5$%.
However, an appropriate relabelling of the pots may bring the FIFA World Cup draw closer to the uniform distribution. Therefore, all possible draw orders of the pots are examined such that the pre-assignment of Russia to Group A is retained. Since the teams can be drawn in $7!\times\left(8!\right)^{3}\approx 3.3\times 10^{17}$ different draw orders, it is almost impossible to derive exact theoretical results. Consequently, the 25 draw mechanisms—the draw procedure of FIFA with the 24 possible draw orders of the pots and the uniformly distributed rejection mechanism (Roberts and Rosenthal,, 2024)—are analysed on the basis of 1 million randomly generated draw orders, analogous to Roberts and Rosenthal, (2024).
The rejection mechanism (or rejection sampler) $U$ works as follows. The groups are drawn completely randomly disregarding any geographical constraint. Then, it is checked whether all restrictions are satisfied. If yes, the draw is retained as a valid one. If no, the current draw is rejected, and a new random group draw is generated. The rejection mechanism is uniform over all possible valid draws, see, for example, Devroye, (1986, Section II.3). In the case of the 2018 FIFA World Cup, about one of 161 random draws is accepted, which is less than one out of 560 as for the 2022 FIFA World Cup (Roberts and Rosenthal,, 2024). Consequently, the total number of valid draws is approximately $7!\times\left(8!\right)^{3}/161\approx 2.05\times 10^{15}$.
@@ 4.1 The effect of the draw order
where $p_{ij}^{M}$ and $p_{ij}^{U}$ are the probabilities that teams $i$ and $j$ are assigned to the same group by a draw mechanism $M$ and the rejection mechanism $U$, respectively. When the biases are aggregated for the teams, we will always sum their absolute value $\lvert\Delta_{ij}\rvert$.
@@
Table 3: Fairness distortions for selected draw orders in the 2018 FIFA World Cup
a Draw order 1-2-3-4
Spain
Peru
Switzerland
England
Colombia
Mexico
Uruguay
Croatia
Denmark
Iceland
Costa Rica
Sweden
Tunisia
Egypt
Senegal
Iran
Serbia
Nigeria
Australia
Japan
Morocco
Panama
South Korea
Saudi Arabia
Russia 3.3 4.1 3.3 3.4 4.2 0.8 4.2 3.3 4.3 4.4 1.1 4.3 2.9 3 2.9 3.1 4.7 2.1 1.7 1.8 2.2 1.9 1.7 1.7
Germany 0.3 0.8 0.4 0.3 0.8 3.7 0.8 0.4 0.5 0.5 0.1 0.5 0.5 0.4 0.4 0.4 1.2 0.5 0 0 0.3 0.2 0.1 0.1
Brazil 2.4 X 2.5 2.4 X 9.8 X 2.5 1 1 1 0.8 0.3 0.4 0.4 0.7 5.2 0.1 0.8 0.9 0 1.6 1 0.9
Portugal 0.4 0.9 0.3 0.4 0.8 3.9 0.9 0.4 0.5 0.5 0.2 0.5 0.4 0.5 0.4 0.3 1.2 0.4 0 0 0.5 0.2 0 0
Argentina 2.4 X 2.5 2.6 X 9.9 X 2.5 1 0.8 0.8 1 0.4 0.4 0.4 0.7 5.2 0 0.9 0.9 0 1.6 0.9 0.9
Belgium 0.2 0.9 0.3 0.3 0.9 3.7 0.9 0.3 0.5 0.6 0.1 0.4 0.5 0.5 0.5 0.2 1.1 0.4 0 0.1 0.5 0.3 0 0
Poland 0.3 0.8 0.4 0.4 0.8 3.8 0.8 0.3 0.4 0.5 0.2 0.5 0.4 0.4 0.4 0.4 1.2 0.4 0 0 0.4 0.3 0.1 0.1
France 0.3 0.8 0.3 0.3 0.9 3.8 0.9 0.3 0.4 0.5 0.1 0.5 0.4 0.3 0.4 0.4 1.1 0.4 0.1 0 0.5 0.2 0 0
Spain X X X X X X X X 0.3 0.4 0.4 0.3 0.3 0.3 0.2 0.6 1.1 0.1 0.1 0.2 0.1 0.2 0.2 0.2
Peru X X X X X X X X 0.6 0.6 0.3 0.7 0.5 0.5 0.5 0.8 2 0.4 0.2 0.4 0.4 0.1 0.4 0.3
Switzerland X X X X X X X X 0.3 0.4 0.3 0.4 0.3 0.3 0.3 0.5 1.1 0.1 0.2 0.2 0.1 0.1 0.2 0.2
England X X X X X X X X 0.4 0.3 0.2 0.4 0.3 0.2 0.2 0.6 1.1 0.1 0.2 0.2 0.1 0.2 0.1 0.2
Colombia X X X X X X X X 0.6 0.6 0.4 0.7 0.5 0.5 0.4 0.8 2 0.5 0.3 0.3 0.5 0.3 0.3 0.4
Mexico X X X X X X X X 0.5 0.4 X 0.4 0.3 0.4 0.4 0.3 10.3 1.7 1.7 1.7 1.7 X 1.7 1.8
Uruguay X X X X X X X X 0.6 0.6 0.4 0.6 0.5 0.5 0.4 0.8 1.9 0.4 0.4 0.2 0.5 0.2 0.3 0.4
Croatia X X X X X X X X 0.4 0.4 0.2 0.3 0.3 0.2 0.2 0.6 1.1 0 0.2 0.2 0.1 0.2 0.2 0.2
Denmark 0.3 0.4 0.4 0.3 0.3 0.3 0.2 0.6 X X X X X X X X 1.7 0.1 0.3 0.3 0.1 0.6 0.3 0.3
Iceland 0.6 0.6 0.3 0.7 0.5 0.5 0.5 0.8 X X X X X X X X 1.7 0 0.2 0.3 0 0.6 0.3 0.2
Costa Rica 0.3 0.4 0.3 0.4 0.3 0.3 0.3 0.5 X X X X X X X X 0.6 0.1 0.2 0.3 0.2 X 0.3 0.2
Sweden 0.4 0.3 0.2 0.4 0.3 0.2 0.2 0.6 X X X X X X X X 1.7 0.1 0.2 0.3 0.1 0.7 0.4 0.3
Tunisia 0.6 0.6 0.4 0.7 0.5 0.5 0.4 0.8 X X X X X X X X 0.7 X 0.2 0.2 X 0.1 0.2 0.2
Egypt 0.5 0.4 X 0.4 0.3 0.4 0.4 0.3 X X X X X X X X 0.8 X 0.2 0.3 X 0.2 0.3 0.1
Senegal 0.6 0.6 0.4 0.6 0.5 0.5 0.4 0.8 X X X X X X X X 0.8 X 0.2 0.2 X 0.1 0.2 0.3
Iran 0.4 0.4 0.2 0.3 0.3 0.2 0.2 0.6 X X X X X X X X 2.2 0 X X 0 2.3 X X
b Draw order 4-2-3-1
Spain
Peru
Switzerland
England
Colombia
Mexico
Uruguay
Croatia
Denmark
Iceland
Costa Rica
Sweden
Tunisia
Egypt
Senegal
Iran
Serbia
Nigeria
Australia
Japan
Morocco
Panama
South Korea
Saudi Arabia
Russia 1.9 2.4 2 2 2.3 0.9 2.4 2 3.6 3.6 5 3.7 1.7 1.8 1.7 0.7 2.1 0 0.5 0.4 0 0.3 0.4 0.4
Germany 0.9 0.5 0.8 0.9 0.4 2.1 0.5 0.9 0.7 0.7 0.2 0.8 0.4 0.6 0.6 0.5 2.5 0 0.5 0.5 0.1 0.4 0.6 0.6
Brazil 1.1 X 1.1 1.2 X 4.6 X 1.1 3.7 3.8 3 3.8 2.2 2.3 2.2 1.7 5.2 0.2 1.1 1.1 0.2 1 1.2 1.2
Portugal 0.8 0.4 0.8 0.8 0.4 2 0.5 0.8 0.8 0.8 0.2 0.8 0.6 0.5 0.5 0.6 2.5 0.1 0.6 0.5 0 0.5 0.5 0.6
Argentina 1.3 X 1 1 X 4.5 X 1.2 3.8 3.8 2.9 3.9 2.3 2.3 2.3 1.7 5.2 0.2 1.1 1.2 0.2 1 1.1 1.1
Belgium 0.9 0.5 0.8 0.8 0.5 2 0.4 0.9 0.8 0.7 0.2 0.8 0.6 0.5 0.5 0.5 2.5 0.1 0.6 0.5 0.1 0.4 0.6 0.6
Poland 0.8 0.4 0.8 0.9 0.5 2 0.5 0.9 0.8 0.8 0.2 0.8 0.6 0.6 0.6 0.5 2.5 0.1 0.5 0.6 0.1 0.5 0.6 0.5
France 0.8 0.5 0.9 0.9 0.6 1.9 0.5 0.8 0.8 0.8 0.2 0.8 0.6 0.6 0.6 0.5 2.5 0.1 0.7 0.6 0 0.4 0.5 0.5
Spain X X X X X X X X 2.5 2.4 1.4 2.6 1.6 1.7 1.6 1.1 3.4 0.6 0.9 1 0.5 0.8 0.9 0.9
Peru X X X X X X X X 2.6 2.5 1.8 2.6 1.6 1.6 1.5 1.3 2.3 0.6 0.6 0.7 0.8 1.1 0.6 0.6
Switzerland X X X X X X X X 2.4 2.5 1.3 2.4 1.6 1.6 1.6 1.2 3.4 0.5 0.9 0.8 0.5 0.9 0.9 0.9
England X X X X X X X X 2.4 2.5 1.4 2.4 1.6 1.6 1.6 1.1 3.4 0.5 0.9 0.9 0.5 0.8 0.9 0.9
Colombia X X X X X X X X 2.6 2.6 1.8 2.6 1.6 1.6 1.5 1.3 2.3 0.7 0.6 0.7 0.7 1.2 0.6 0.6
Mexico X X X X X X X X 2.1 2.1 X 2 1.7 1.8 1.8 0.8 6.8 0 1.7 1.6 0.1 X 1.6 1.7
Uruguay X X X X X X X X 2.6 2.6 1.9 2.6 1.5 1.6 1.6 1.2 2.2 0.7 0.7 0.6 0.6 1.1 0.6 0.6
Croatia X X X X X X X X 2.5 2.4 1.4 2.5 1.7 1.6 1.7 1.1 3.4 0.5 0.9 0.8 0.5 0.8 0.8 0.9
Denmark 2.5 2.4 1.4 2.6 1.6 1.7 1.6 1.1 X X X X X X X X 0.2 0.2 0.1 0.2 0.2 0.3 0 0.1
Iceland 2.6 2.5 1.8 2.6 1.6 1.6 1.5 1.3 X X X X X X X X 0.2 0.3 0.1 0 0.2 0.3 0.2 0.1
Costa Rica 2.4 2.5 1.3 2.4 1.6 1.6 1.6 1.2 X X X X X X X X 0.7 0.9 0.7 0.5 0.9 X 0.6 0.7
Sweden 2.4 2.5 1.4 2.4 1.6 1.6 1.6 1.1 X X X X X X X X 0.2 0.2 0.2 0.1 0.2 0.3 0 0.1
Tunisia 2.6 2.6 1.8 2.6 1.6 1.6 1.5 1.3 X X X X X X X X 0.1 X 0.2 0 X 0.3 0.1 0.1
Egypt 2.1 2.1 X 2 1.7 1.8 1.8 0.8 X X X X X X X X 0.1 X 0.1 0.1 X 0.4 0.2 0.2
Senegal 2.6 2.6 1.9 2.6 1.5 1.6 1.6 1.2 X X X X X X X X 0.1 X 0.1 0.1 X 0.3 0.1 0.1
Iran 2.5 2.4 1.4 2.5 1.7 1.6 1.7 1.1 X X X X X X X X 0.3 0.2 X X 0.3 0.1 X X
  • X represents a pair of teams that cannot play in the group stage.
  • The numbers show percentages ($100\times\Delta_{ij}$) rounded to one decimal place.
  • Green (Red) colour means that the FIFA draw procedure implies a higher (lower) probability than a uniform draw. For instance, the probability of assigning Mexico and Serbia to the same group is higher (lower) by 10.3 (6.8) percentage points under the FIFA draw mechanism with the draw order 1-2-3-4 (draw order 4-2-3-1) than under a uniform draw when it is 22.7%.
  • Darker colour indicates a higher value.
Table 3 presents fairness distortions for two fundamentally different draw orders. Clearly, Russia and Mexico pose challenges for both of them. However, the official mechanism (draw order 1-2-3-4) is clearly better for the country pairs from Pot 1 and Pot 2, as well as from Pot 1 and Pot 3, which is favourable as these teams have the highest chance to qualify for the Round of 16.
@@
The distortions are also worth studying by taking the average of absolute or squared biases for all country pairs that involve a given national team. For the sake of simplicity, four draw orders are analysed in detail: the official 1-2-3-4, one that improves both the minimal nonzero and maximal probabilities with a slight change (1-2-4-3), one that is less biased for Mexico (1-3-2-4), and a fundamentally different draw order (4-2-3-1).
@@
Figure 3 compares the means of the biases for the 32 countries. The official draw procedure is far from the fair rejection mechanism $U$ in the case of Russia (1), Brazil (3), Argentina (5), Mexico (14), and Serbia (25). Note that Brazil and Argentina are interchangeable in the draw, thus, the deviations for these two countries are different only due to the (seemingly low) sampling error. Reversing the draw order of Pots 3 and 4 reduces unfairness for Russia and Serbia but it does not treat the problem for Mexico. Contrarily, the deviation for Costa Rica (19) strongly increases with the draw order 1-2-4-3. Drawing Pot 3 immediately after the best teams in Pot 1 implies a relatively low distortion in the case of Russia, and, especially, Mexico but the situation of the African (21, 22, 23) and European (17, 18, 20) teams in Pot 3 becomes less fair compared to the draw order 1-2-4-3. Finally, the draw order 4-2-3-1 is unfair with respect to all countries in Pot 2 except for Mexico.